nt Chapter 21, Problem 047 Point charges of q1-6.08 uC and 2451 C are placed on an xas, a x.15 m and x-15.2m, respetivly, a h must be placed at x 21.6 m so that any charge q4 placed at the origin would experience no electrostatic force? n theeare betm Whia Sharge on p nu
Answer #1

As given that net electrostatic force on q4 = 0

Force on q4 due to other charges,

kq1q4 / x1^2 + kq2q4 / x2^2 + kq3q4 / x3^2 = 0

k*q4 (q1 / x1^2 + q2 / x2^2 + q3 / x3^2) = 0

(q1 / x1^2 + q2 / x2^2 + q3 / x3^2) = 0

(6.08*10-6 / 7.152) - (4.51*10-6 / 15.22) + (q3 / 21.62) = 0

q3 = -46.38*10-6 C

q3 = -46.38 uC